# Computer Vision Datasets and Models Exhibit Cultural and Linguistic Diversity in Perception

### v3

@@ 6.2 Limitations
Risk of cultural essentialism. Categorizing differences solely with languages may pose a risk of essentializing or stereotyping cultures, suggesting that all members that speak a language perceive the world similarly.
6.3 Recommendations
Dataset construction. Most general image datasets do not provide background information about annotators. Disregarding the cultural background of annotators can unduly overrepresent a set of perceptual behaviors over others. Datasets should also collect annotator background, such as their country or culture, their primary language, and other psychological factors that might substantively affect an individual’s perception. At the minimum, unique annotator identifiers should be included to allow linking of different annotations to their annotator. This information is generally cheap to collect, but will significantly aid the development of perception-aware models and applications.
Multilingual modeling. Currently, many multilingual vision-language models are built by training models on captions translated from English into the desired language. Whenever possible, multilingual vision-language applications should be built by training models collected from native speakers. Our results suggest that the advantage is not only syntactic (i.e. more “natural-sounding” than translations), but that annotations produced in the desired language describe different semantic content and manners of expression.
Multilinguality as perceptual diversity. Multilingual vision-language datasets should be explored as repositories of perceptually diverse annotations. At a basic level, simply training models on multilingual rather than monolingual vision datasets may expose them to a more diverse, rich set of concepts. At a more complex level, specifically designed structured learning paradigms such as contrastive learning may be able to exploit concept diversity across multilingual annotations to yield richer image representations.
Designing for diverse perception. Vision-language applications, such as image captioning, image search, and image generation, should consider, or even condition results on, users’ perceptual background. New systems or models should report metrics across culturally diverse tests sets, and consider running human evaluations with users from culturally diverse backgrounds.
6.4 Conclusion
Similar to evidence presented in social science studies, we provide empirical evidence for cultural and linguistic differences in perception within computer vision datasets and models. Methodologically, we turn the focus from an objective conception of the “ground truth” towards the human subject, which selectively attends to certain objects and qualities in the nearly-infinite field of raw visual features [77, 73]. Our work emphasizes the need for computer vision to understand the cultural [80] and linguistic [64] dimensions of this human subject.
Appendix A Experimental Details
A.1 Translation into English
@@
| More appropriate to use ”memories” instead of ”impressions” |
@@
| Particle suggests that numbers are written “using” sheet of paper, not “on” it. |
@@
| “toile d’ araignée” can be directly translated to “cobweb” |
A.2 Probing Multilingual Capabilities in LLaVA
@@
| Roles | English |
(user, assistant) |
| French |
(utilisateur, assistant) |
| German |
(Benutzer, Assistent) |
| System | English |
A conversation between a user and an LLM-based AI assistant. The assistant gives helpful and honest answers. |
| French |
Une conversation entre un utilisateur et un assistant IA basé sur LLM. L’assistant donne des réponses utiles et honnêtes. |
| German |
Ein Gespräch zwischen einem Benutzer und einem auf LLM basierenden KI-Assistenten. Der Assistent gibt hilfreiche und ehrliche Antworten. |
| User Prompt | English |
What is in this image? Answer in English. |
| French |
Qu’est-ce qu’il y a dans cette image? Répondez en français. |
| German |
Was ist auf diesem Bild? Antwort auf Deutsch. |
A.3 Image Captioning User Study
We recruited 10 English speakers from the US and 10 Japanese speakers from Japan (referenced in follow-up to semantic content evaluation). The instructions given to them are presented in Figures 2(a) and 2(b). A sample of the produced captions is given in Table 10.
English | I. |
Two very small boats on a river |
| II. |
Toy boats in the water |
| III. |
A yellow boat and a red boat that appear to be models. |
@@ A.3 Image Captioning User Study
Two boats on the water and a building in the back |
| II. |
close-up of a model of a boat and people on the waterfront |
| III. |
Two boats floating on the river and a model of the town in the distance |
English | I. |
Luggage left unattended at a table. |
| II. |
Luggage lined up next to tables with jackets resting on the tables. |
| III. |
luggage sitting next to tables |
@@
A man sitting in a lobby with lots of suitcases and bags |
| II. |
A man is sitting in a room, and there are several tables filled with luggage nearby. |
| III. |
Man waiting with a lot of luggage |
English | I. |
Cat laying down in an arm chair. |
| II. |
A Siamese cat laying on its back on a couch next to a pillow. |
| III. |
A cat stretched out and upside down on a chair |
@@
A cat stretches out on a blue chair and a pillow with an embroidered owl next to it. |
| II. |
A cat is relaxing next to a cushion with a picture of an owl on it. |
| III. |
Cat sitting on his back in an armchair with an owl-patterned cushion |
Appendix B Supplementary Data and Figures
Our results across all evaluations are displayed in Table 11.
B.1 Semantics Evaluations
Figure 4 shows that despite an expected diminishing-returns trajectory, continuously unioning even a well-developed existing scene graph with a new language’s scene graph expands it. This suggests that different languages continue to have new information to add to the existing scene graph of visual knowledge. Table 12 displays the sizes of intersections between monolingual scene graphs as measured by the number of objects and relations, using the formula . It is an alternative way to understand the conceptual overlap of different languages. Table 13 shows that scene graphs constructed from captions from the same model but different languages are only slightly smaller than those constructed from captions from the same language but different models. (Referenced in follow-up .) Table 14 shows the intersection sizes between monolingual and multi-model scene graphs. (Referenced in follow-up .) Table 3 shows the sizes of concept sets under intersection with objects documented in Visual Genome. (Referenced in follow-up .) Table 15 shows some samples in which multilingual scene graphs identify objects in the image which are not mentioned in the Visual Genome annotations. (Referenced in follow-up .) Figure 5 shows several examples of scene graphs generated in different languages for different samples.
B.2 Model outputs evaluations
Tables 16 and 17 repeat the same fine-tuning experiment as outlined in the model generative outputs evaluation, but training on LLaVA and XM captions instead of Vertex captions.
@@ B.2 Model outputs evaluations
| en | fr | de | ru | zh | ja | ko | en,fr,zh | fr,zh,ru | de,fr,ru | avg222The ‘avg’ column is computed by randomly sampling across all languages. | ||||
@@
| LLaVA | Objects | 4.54 | 5.05 | 5.26 | 4.52 | 4.54 | - | 333Japanese and Korean are excluded from LLaVA results due to extremely poor caption generation in the languages. - | 6.14 | 6.15 | 6.25 | 5.93 | ||
@@
| Fine-tuning | Vertex | Best Model444Each cell value is the original language of the caption set used to train the model which performed best as evaluated by the caption set originally from the language specified in the column. | en | de | fr | ru | zh | ja | ko | - | - | - | multi | |
@@
| Language | ||||||||
| en | de | fr | ru | zh | ja | ko | mm | |
| Objects | 1.96 | 1.92 | 1.93 | 1.97 | 1.85 | 1.73 | 1.76 | 3.59 |
| Relations | 0.79 | 0.76 | 0.74 | 0.78 | 0.70 | 0.62 | 0.64 | 2.51 |
@@
giraffe tail, spot, rock, giraffe, rocks, grass |
bird (left of image) |
| Evaluated on | ||||||
| en | de | fr | ru | zh | multi | |
| en | 0.225 | 0.229 | 0.219 | 0.218 | 0.230 | |
| de | 0.213 | 0.219 | 0.217 | 0.215 | 0.219 | |
| fr | 0.248 | 0.240 | 0.234 | 0.236 | 0.246 | |
| ru | 0.226 | 0.234 | 0.228 | 0.231 | 0.239 | |
| zh | 0.199 | 0.202 | 0.199 | 0.207 | 0.216 | |
| multi | ||||||
| Evaluated on | ||||||||
| en | de | fr | ru | zh | ja | ko | multi | |
| en | 0.124 | 0.1421 | 0.120 | 0.114 | 0.129 | 0.130 | 0.148 | |
| de | 0.158 | 0.152 | 0.143 | 0.124 | 0.140 | 0.146 | 0.149 | |
| fr | 0.182 | 0.142 | 0.143 | 0.130 | 0.146 | 0.150 | 0.154 | |
| ru | 0.172 | 0.136 | 0.152 | 0.125 | 0.137 | 0.142 | 0.148 | |
| zh | 0.144 | 0.116 | 0.129 | 0.120 | 0.130 | 0.142 | 0.130 | |
| ja | 0.144 | 0.128 | 0.137 | 0.125 | 0.124 | 0.144 | 0.135 | |
| ko | 0.151 | 0.116 | 0.131 | 0.116 | 0.115 | 0.134 | 0.134 | |
| multi | ||||||||
![[Uncaptioned image]](extracted/5459933/imgs/user-study-example-panel/example_caption_image_user_study2.png)
![[Uncaptioned image]](extracted/5459933/imgs/user-study-example-panel/example_caption_image_user_study3.png)
![[Uncaptioned image]](extracted/5459933/imgs/user-study-example-panel/example_caption_image_user_study4.png)
![[Uncaptioned image]](extracted/5459933/imgs/vg-exceed-panel/vg_example8.png)
![[Uncaptioned image]](extracted/5459933/imgs/vg-exceed-panel/vg_example2.png)
![[Uncaptioned image]](extracted/5459933/imgs/vg-exceed-panel/vg_example3.png)
![[Uncaptioned image]](extracted/5459933/imgs/vg-exceed-panel/vg_example4.png)
![[Uncaptioned image]](extracted/5459933/imgs/vg-exceed-panel/vg_example7.png)

### v4

@@ 7 Limitations
Risk of linguistic essentialism. Categorizing differences solely with languages may pose a risk of essentializing or stereotyping them, suggesting that all members that speak a language describe the world similarly. We emphasize that we do not make categorical but rather distributional claims, aiming to show general differences across a large set of samples.
8 Experimental Details
8.1 Translation into English
@@
One possible confounder in results like §4 is that language-specific syntactic artifacts introduced during translation. For instance, text translated from German into English might have a unique syntactic structure which distinguishes it from text originally written in English. If this is the case, then it should be possible to identify translated text from one language versus another. To test this limitation, we embed all translated captions using a BERT-based model [100]. We fit a logistic regression model to predict a sample’s original language from these features, and find near-random chance performance at (random chance is ). This suggests that the translation artifact confounder does not explain the observed results.
8.2 Probing Multilingual Capabilities in LLaVA
8.3 Image Captioning User Study
We recruited 10 English speakers from the US and 10 Japanese speakers from Japan. The instructions given to them are presented in Figures 3(a) and 3(b). A sample of the produced captions is given in Table 9.3.
Because large-scale image-text datasets do not conduct much annotator information, it is difficult to make detailed and strong inferences about the psychological causes of the observed results, so more work is needed in this direction. However, as a start, we recruited 10 English speakers from the U.S. and 10 Japanese speakers from Japan to caption 30 Visual Genome images and repeated the semantic content evaluation for human-produced captions. We find, in the same pattern as before with model captions, that unioning English scene graphs with Japanese scene graphs expands the size by objects, relations, and attributes over unioning English scene graphs with other English scene graphs. Moreover, a manual inspection of the captions suggests that the captions roughly echo the predictions from cross-cultural perceptual psychology – Japanese captions tend to mention background objects and information more than English ones (see Figures 1 and 9.3).
9 Supplementary Data and Figures
Our results across all evaluations are displayed in Table 10.
9.1 Semantics Evaluations
Figure 9.3 shows that despite an expected diminishing-returns trajectory, continuously unioning even a well-developed existing scene graph with a new language’s scene graph expands it. This suggests that different languages continue to have new information to add to the existing scene graph of visual knowledge. Table 11 displays the sizes of intersections between monolingual scene graphs as measured by the number of objects and relations, using the formula . It is an alternative way to understand the conceptual overlap of different languages. Table 12 shows that scene graphs constructed from captions from the same model but different languages are only slightly smaller than those constructed from captions from the same language but different models. Table 13 shows the intersection sizes between monolingual and multi-model scene graphs. Table 16 shows some samples in which multilingual scene graphs identify objects in the image which are not mentioned in the Visual Genome annotations. Figure 5 shows several examples of scene graphs generated in different languages for different samples.
9.2 Multilingual Embedding Space Coverage
Recall from §2.1 that many of the tools we use to measure semantics and expressions are not available in different languages (e.g., scene graph parsers, linguistic measures). However, in the case of embedding space coverage, we can use multilingual embeddings rather than monolingual (English) embeddings (with translation of all captions into English). We reproduce the expressive variation experiment described in §2.3 using multilingual embeddings without translation, and find that the same result holds as in the main paper using English embeddings with translation 5. This provides further empirical support that translation bias does not interfere with our results. However, note that multilingual embeddings have documented language biases [87, 18], which is why we prefer to use monolingual embeddings with translation for a fairer comparison.
| mono | multi | |||||
| en | fr | avg | en,fr,de | en,ru,zh | avg | |
| XM | .274 | .279 | .280 | .327 | .328 | .340 |
| LLaVA | .475 | .507 | .521 | .704 | .795 | .753 |
| Vertex | .340 | .321 | .321 | .600 | .647 | .612 |
9.3 Model outputs evaluations
Tables 14 and 15 repeat the same fine-tuning experiment as outlined in §4, but training on LLaVA and XM captions instead of Vertex captions.
@@ 9.3 Model outputs evaluations
| More appropriate to use "memories” instead of "impressions" |
@@
| Particle suggests that numbers are written “using" sheet of paper, not “on" it. |
@@
| “toile d’ araignée" can be directly translated to “cobweb" |
@@
| Roles | English | (user, assistant) |
| French | (utilisateur, assistant) | |
| German | (Benutzer, Assistent) | |
| System | English | A conversation between a user and an LLM-based AI assistant. The assistant gives helpful and honest answers. |
| French | Une conversation entre un utilisateur et un assistant IA basé sur LLM. L’assistant donne des réponses utiles et honnêtes. | |
| German | Ein Gespräch zwischen einem Benutzer und einem auf LLM basierenden KI-Assistenten. Der Assistent gibt hilfreiche und ehrliche Antworten. | |
| User Prompt | English | What is in this image? Answer in English. |
| French | Qu’est-ce qu’il y a dans cette image? Répondez en français. | |
| German | Was ist auf diesem Bild? Antwort auf Deutsch. | |
English | I. | Two very small boats on a river |
| II. | Toy boats in the water | ||
| III. | A yellow boat and a red boat that appear to be models. | ||
@@
| II. | close-up of a model of a boat and people on the waterfront | ||
| III. | Two boats floating on the river and a model of the town in the distance | ||
English | I. | Luggage left unattended at a table. |
| II. | Luggage lined up next to tables with jackets resting on the tables. | ||
| III. | luggage sitting next to tables | ||
@@
| II. | A man is sitting in a room, and there are several tables filled with luggage nearby. | ||
| III. | Man waiting with a lot of luggage | ||
English | I. | Cat laying down in an arm chair. |
| II. | A Siamese cat laying on its back on a couch next to a pillow. | ||
| III. | A cat stretched out and upside down on a chair | ||
@@
| II. | A cat is relaxing next to a cushion with a picture of an owl on it. | ||
| III. | Cat sitting on his back in an armchair with an owl-patternedcushion |
@@
| en | fr | de | ru | zh | ja | ko | en,fr,zh | fr,zh,ru | de,fr,ru | avg111The ‘avg’ column is computed by randomly sampling across all languages. | ||||
@@
| LLaVA | Objects | 4.54 | 5.05 | 5.26 | 4.52 | 4.54 | - | 222Japanese and Korean are excluded from LLaVA results due to extremely poor caption generation in the languages. - | 6.14 | 6.15 | 6.25 | 5.93 | ||
@@
| Fine-tuning | Vertex | Best Model333Each cell value is the original language of the caption set used to train the model which performed best as evaluated by the caption set originally from the language specified in the column. | en | de | fr | ru | zh | ja | ko | - | - | - | multi | |
@@
| Language | |||||||||
| en | de | fr | ru | zh | ja | ko | mm | ||
| Objects | 1.96 | 1.92 | 1.93 | 1.97 | 1.85 | 1.73 | 1.76 | 3.59 | |
| Relations | 0.79 | 0.76 | 0.74 | 0.78 | 0.70 | 0.62 | 0.64 | 2.51 | |
Metric |
@@
| Evaluated on | |||||||
| en | de | fr | ru | zh | multi | ||
| en | 0.225 | 0.229 | 0.219 | 0.218 | 0.230 | ||
| de | 0.213 | 0.219 | 0.217 | 0.215 | 0.219 | ||
| fr | 0.248 | 0.240 | 0.234 | 0.236 | 0.246 | ||
| ru | 0.226 | 0.234 | 0.228 | 0.231 | 0.239 | ||
Fine-tuned on |
zh | 0.199 | 0.202 | 0.199 | 0.207 | 0.216 | |
| multi | |||||||
| Evaluated on | |||||||||
| en | de | fr | ru | zh | ja | ko | multi | ||
| en | 0.124 | 0.1421 | 0.120 | 0.114 | 0.129 | 0.130 | 0.148 | ||
| de | 0.158 | 0.152 | 0.143 | 0.124 | 0.140 | 0.146 | 0.149 | ||
| fr | 0.182 | 0.142 | 0.143 | 0.130 | 0.146 | 0.150 | 0.154 | ||
| ru | 0.172 | 0.136 | 0.152 | 0.125 | 0.137 | 0.142 | 0.148 | ||
| zh | 0.144 | 0.116 | 0.129 | 0.120 | 0.130 | 0.142 | 0.130 | ||
Fine-tuned on |
ja | 0.144 | 0.128 | 0.137 | 0.125 | 0.124 | 0.144 | 0.135 | |
| ko | 0.151 | 0.116 | 0.131 | 0.116 | 0.115 | 0.134 | 0.134 | ||
| multi | |||||||||
@@
giraffe tail, spot, rock, giraffe, rocks, grass | bird (left of image) |
![[Uncaptioned image]](extracted/6297941/imgs/user-study-example-panel/example_caption_image_user_study2.png)
![[Uncaptioned image]](extracted/6297941/imgs/user-study-example-panel/example_caption_image_user_study3.png)
![[Uncaptioned image]](extracted/6297941/imgs/user-study-example-panel/example_caption_image_user_study4.png)
![[Uncaptioned image]](extracted/6297941/imgs/vg-exceed-panel/vg_example8.png)
![[Uncaptioned image]](extracted/6297941/imgs/vg-exceed-panel/vg_example2.png)
![[Uncaptioned image]](extracted/6297941/imgs/vg-exceed-panel/vg_example3.png)
![[Uncaptioned image]](extracted/6297941/imgs/vg-exceed-panel/vg_example4.png)
![[Uncaptioned image]](extracted/6297941/imgs/vg-exceed-panel/vg_example7.png)